Paige is one fifth her mothers age if the sum of their age is 30 how old is paige?
Answer by hamsanash1981@gmail.com(151)

Let Paige's mother age be x and Paige age = (1/5)X = > x/5
sum of their age is 30
=> x + x/5 = 30
(5x + x)/5 = 30
6x = 30*5
x = 25
Paige age = x/5 => 25/5
Paige's age = 5